Comupute all the minors and the cofactors of: [1 2 3]

[2 0 1]

[2 3 4]


1
Expert's answer
2021-05-24T15:29:55-0400

Solution.


"A=\begin{pmatrix}
 1& 2&3\\
 2& 0&1\\
2&3&4
\end{pmatrix}"

The elements of the minor of the matrix A will be given by

"M_{1,1}=\begin{vmatrix}
 0 & 1\\
 3& 4
\end{vmatrix}=-3"

"M_{1,2}=\begin{vmatrix}
 2& 1\\
 2& 4
\end{vmatrix}=6"

"M_{1,3}=\begin{vmatrix}
 2 & 0\\
 2& 3
\end{vmatrix}=6"

"M_{2,1}=\begin{vmatrix}
 2& 3\\
 3& 4
\end{vmatrix}=-1"

"M_{2,2}=\begin{vmatrix}
 1 & 3\\
 2& 4
\end{vmatrix}=-2"

"M_{2,3}=\begin{vmatrix}
 1 & 2\\
 2& 3
\end{vmatrix}=-1"

"M_{3,1}=\begin{vmatrix}
 2 & 3\\
 0& 1
\end{vmatrix}=2"

"M_{3,2}=\begin{vmatrix}
 1& 3\\
 2& 1
\end{vmatrix}=-5"

"M_{3,3}=\begin{vmatrix}
 1& 2\\
 2& 0
\end{vmatrix}=-4"

Hence, "M=\begin{pmatrix}
 -3& 6&6 \\
 -1& -2&-1\\
2&-5&-4
\end{pmatrix}"

The cofactor matrix will be given by "C_{i,j}=(-1)^{i+j}M_{i,j}."

"C=\begin{pmatrix}
 -3& -6&6 \\
 1& -2&1\\
2&5&-4
\end{pmatrix}"